Suppose 1000 units of a product A are produced by XYZ limited but the quantity demanded is 2000 units.A $18 change in the price of product A results in a change in quantity demanded and supplied of 6 and 9 units.XYZ has a workforce of 500 people who pay income tax of $200 to the government.
A.Calculate the equilibrium price and equilibrium quantity.
B. Suppose the government introduces a subsidy of $10 on each unit of product A produced. Calculate the new equilibrium price and quantity.
Graphically indicate how much of the subsidy will be received by the producer and consumer .
1
Expert's answer
2020-03-05T08:57:53-0500

A. Using the equation of a line:

1000 + 9x = 2000 - 6x,

15x = 1000,

x = 66.67,

Qd = Qs = 1000 + 9*66.67 = 1600 units.

Equilirium price can't be determined without additional information provided.

B. If the government introduces a subsidy of $10 on each unit of product A produced, then the new equilibrium price will decrease and quantity will increase. 

There is not enough information to calculate how much of the subsidy will be received by the producer and consumer.